Que- tan(A/2) is equal to
1)cosecA + cotA
2)secA - cotA
3)cosecA - cotA
4)secA + cotA

Grade:12th pass

1 Answers

sahil
6 years ago
Tan (a/2)=sin(a/2)/cos(a/2)=2sin^2(a/2)/2sin(a/2)cos(a/2)=1-cos(a)/sin(a)=cosec(a)-cot(a) option 3 is correct .Please approve if it helps